# Exploring Value Biases: How LLMs Deviate Towards the Ideal

### 16 Feb 2024 v1

@@ 1 Introduction
Large language models (LLMs) are being deployed across a wide range of applications (Brown et al., 2020; Schick & Schütze, 2021) with an increasing socio-economic impact (Eloundou et al., 2023). The quality of responses generated by LLMs is an essential area of research from both security (Yao et al., 2023b) and utility perspectives (Achiam et al., 2023). LLMs have been found to be capable of doing mathematical reasoning (Imani et al., 2023), pragmatics (Lipkin et al., 2023), and creating high-level plans (Song et al., 2023) to varying degrees even without specific deliberation methods like CoT (Wei et al., 2022). Understanding this non-deliberate mechanism of efficient response sampling is crucial to understanding its performance and biases.
Previous work evaluating the biases in the response of LLMs has shown them to have significant gender-bias (Kotek et al., 2023), social-stereotypes (Kotek et al., 2023), and popularity bias (Hou et al., 2023). Studies have also been made specific to critical applications like medical domain (Omiye et al., 2023), code generation (Huang et al., 2023), and to biases specific to cultures (Khandelwal et al., 2023) or geographic regions (Manvi et al., 2024). These are relevant studies for real-world applications but often have context-based confounders. It is essential to focus on understanding the underlying foundational biases manifesting in different forms of bias.
In this work, we study the sampling of LLM in light of value bias and show that it tends to favour high-value options across different scenarios. Since the underlying auto-regressive mechanism is not goal-driven, it is not trivial how, from a large number of possibilities of samples, LLM chooses based on value. This value bias is a tendency for the sample to deviate from the (statistically) high likely value towards a notion of ‘ideal value’. This unexpected bias can manifest as different biases based on the value system and the evaluated context.
@@
We show that value bias manifests in cases where the notion of ‘ideal’ stems from an implicit value system of the LLM or a value system learnt in context. We motivate our experiments from the ideas presented by Phillips et al. (2019), which states possibility sampling comes from a mixture of likelihood and value. We use psychology-inspired experiments for value-bias learnt in context (Bear et al., 2020), implicit value-bias (Bear et al., 2020) and show the manifestation of value bias in forming normality and prototypes (Bear & Knobe, 2017a). The notion of ‘ideal’ can manifest in other applications like text summarization, reducing response quality.
Value bias can be useful in many applications, given the value system of the LLM aligns with the use case. Under the categories evaluated, our experiment shows a lack of correlation between the value system of humans and LLM. Our result shows how LLMs can learn a value system in context with a single prompt with a statistically significant shift of the sample towards the ideal side of the average, as illustrated in Figure 1.
@@
  • We show that the LLM has an implicit value bias for concepts. We also show that value bias manifests even for new concepts learnt in context (zero-shot).
  • We show the practical implications of value bias in LLMs’ judgment of concept prototypicality. We also show that the value bias in LLM might not align with the value system of humans.
@@ 2.1 LLM Biases
Previous work on LLM predominantly evaluates biases with respect to social concepts like gender, race, and popularity. There has also been investigation of biases in aspects like language style and lexical content (Wan et al., 2023). Gallegos et al. (2023) comprehensively surveys these works and presents a taxonomy of biases. This taxonomy aligns with how humans attribute meanings to these biases and their impact on society. The biases of LLMs have also been studied in the context of specific fields and applications like health care  (Omiye et al., 2023; Zack et al., 2024; Thirunavukarasu et al., 2023; Hastings, 2024). These studies, however, do not go beyond the human taxonomy of biases to explore fundamental biases that, in turn, manifest in real-world applications.
@@ 2.2 Understanding Heuristics in Response Sampling
Human reasoning has been broadly characterised as a two-step process of quick ‘System-1’ reasoning and a more deliberate ‘System-2’ reasoning (Daniel, 2017). Interestingly, LLMs have been compared to System1 thinking due to their automatic and quick nature (Yao et al., 2023a). In fact, recent work shows some overlaps in errors made by LLMs and humans in System1 reasoning tasks (Dasgupta et al., 2022).
Simon (1996) uses the notion of heuristics in the decision-making of ‘System-1’ mechanisms. They show the utility of ‘mental shortcuts’ to navigate countless possibilities of the search problem (Newell et al., 1972). Exploring the heuristics of LLMs can help understand the information processing in them (Hazra et al., 2023; Shah et al., 2023; Suri et al., 2023). The mechanism by which LLMs sample their response and the heuristics followed require further investigation.
System1 outputs play a considerable role in System2 processing because the latter depends on the former as a prior for decision making. For instance, in AlphaGo (Silver et al., 2016), the deliberate system, i.e. the MCTS algorithm, relies on the estimates of the neural network to cut down the search space of solutions. Similarly, sampling of the LLM becomes critical for the deliberate problem solver to narrow down the space of possibilities. For example, in ToT (Yao et al., 2023a), when the LLM is combined with a symbolic solver, there is an inherent assumption that LLMs would give a good prior for the symbolic system to solve for it to execute accurate deliberation. Even for problem-solving at ‘System-2’, understanding sampling heuristics is of much significance.
@@ 2.3 Possibility Sampling in Humans
A growing body of literature addresses how humans and animals navigate the search problem of countless possibilities (Phillips et al., 2019; Phillips & Cushman, 2017; Mattar & Lengyel, 2022; Ross et al., 2023). In this line of work, human decision-making is thought of as a two-step process where model-free reinforcement learning is looked at as a cheap and efficient way to construct a consideration set of possible options and spend computational resources in a model-based framework to arrive at the final choice among the limited sampled candidates (Gershman et al., 2014). Increasing evidence shows the significance of likelihood and value in how humans sample possible solutions from the large space of possibilities Bear et al. (2020); Phillips et al. (2019); Bear & Knobe (2017b).
In short, what comes to the human mind or what it considers normal is sampled from a probability distribution of both value and probability (frequency) (Bear et al., 2020). This dual nature of thought is hypothesised to have come due to humans being goal-driven agents and engaging in value maximisation (Bear & Knobe, 2017b). In spatial navigation, experiments show that an optimal reinforcement learning agent which has the ability to recall past experiences, ordered on utility, optimises on two dimensions, namely, "gain" and "need", where gain roughly refers to the value aspect of experience while need refers to the statistical occurrence of experience Mattar & Daw (2018).
Other efforts try to encode the possibility and value dimensions into the modalities ’could’ and ’should’ where ’could’ shows how humans imagine possible alternatives, and ’should’ shows the high-value possibilities. Poulsen  et al. show the influence of modal value when the LLM consider possibilities Poulsen & DeDeo (2023).
Figure 2: The sample of an LLM drifts away from the average value towards an ideal value. The notion of the ideal is associated with the category we evaluate. For some prompts, the ideal value is greater than the average value and sometimes the other way, depending on the value system associated with the concept. $\alpha$ shows the deviation in the direction of the ideal. alpha shows the positive value bias.
@@ 3.1 Implicit Bias
We evaluate the likelihood of samples and the notion of the ideal on $n$ different categories $C^{1-n}$. The average value reported by the LLM for a category $C$ is ${C}_{a}$ and the ideal value is ${C}_{i}$. The sample chosen by the LLM for the category is $C_{s}$. We use a variable $\alpha$ to quantify the degree of value bias. For a positive value bias, the $C_{s}$ deviate from the $C_{a}$ in the direction of $C_{i}$. We compute this direction as the positive direction of alpha.
@@
The sign of alpha is given by the relative positions of $C_{a}$,$C_{i}$ and $C_{s}$ such that $C_{s}$ is in the direction of $C_{i}$ as measured from $C_{a}$. Figure 2 shows few different configurations of $C_{a}$,$C_{i}$ and $C_{s}$ and the corresponding $\alpha$.
We also compute $\alpha$, a normalized scale such that $C_{a}$ is at the origin and $C_{i}$ is at unit distance from the origin in the positive direction. We call this $\hat{\alpha}$, and it enables comparison across categories with less dependency on the scale of values. It also allows comparison with values obtained in human experiments. Value bias is significant with a higher positive value of $\hat{\alpha}$. We compute $\hat{\alpha}$ as
@@ 3.2 Evaluating value bias for in-context value and distributions
For the new category $C$ and the corresponding $N$ and $V$ introduced in context, we ask the LLM to report the average of the distribution $C_{a}$ and a sample from the distribution $C_{s}$. Note that the notion of ideal comes from $V$. To test the significance of value bias, we use the Mann-Whitney-U test to compare $C_{a}$ to $N$ and $C_{a}$ to $C_{s}$ given $V$.
@@ 3.3 Value bias in evaluating prototypes
To evaluate the potential effect of value bias in applications, we assess their performance in evaluating prototypes. Prototypes are a means of understanding how people categorize objects and concepts. They are often considered the “best example” or an “average” representation of a category. For instance, when thinking of a category like “birds”, the prototype would be a common bird like a sparrow or a robin, which most closely matches the average characteristics of the category Smith & Medin (1981).
We evaluate if the judgement of LLM on the prototype is driven solely by the statistical likelihood or biased by the value component. We provide a category $C$ and a corresponding exemplar. We ask the LLM to judge on three dimensions, namely the average ($C_{a}$), ideal ($C_{i}$) and prototypicality ($C_{p}$) of the exemplar. We test for significance with a binomial test if the sampled prototypicality $C_{p}$ is pushed away from $C_{a}$ in the direction of $C_{i}$.
@@ 4 Experiments and Results
In this section, we present the three experiments for showing value bias: (a) implicit with each concept, (b) learnt in context and (c) in prototypes of concepts. Our results show a significant value bias in the sampling of LLM. Unless mentioned otherwise, we use GPT-4 for all our experiments. Also, all prompts were appended with a pre-prompt to get floating-point numbers as answers. The complete text used in the prompts for each experiment are given in the Appendix.
@@ 4.1 Implicit value bias
We use a third prompt to get a sample $C_{s}$ from each category. To get a value of $C_{s}$, we use a prompt similar to the question used in human study Bear et al. (2020). For instance, the prompt for the category ‘TV watching of people’ is as follows:
$P_{s}$: Enter the first number that comes to mind when answering the question. Note that there is no correct answer. Please be spontaneous in the judgement: What is the number of hours of TV for a person to watch in a day?
The value reported by the LLM is averaged over ten runs reported in Table  1. The entries where $\alpha$ is positive are shown in purple colour. These categories show a value bias in sampling.
@@
Table 1: The table shows the average, ideal and sample values for the 36 different categories. The categories with a positive $\alpha$ (showing a value bias) are shown in purple colour.
Of the 1200 runs, 21 experiments were rejected due to the LLM not generating scores for the prompt. 1200 runs correspond to the ten iterations of the three prompts across 40 categories: average, ideal and sample. We hypothesize that the LLM might need a jailbreak to give output to queries like: “Ideal amount of parking tickets a person should get in a year”. We analyze the 36 categories where we get consistent results.
From Table 1, we can see that 24 out of 36 categories are on the ideal side of the average when sampled ($p$ = 0.033), which is statistically significant, indicating a value bias in sampling. The sample is expected to lie in between the average and ideal due to the value bias. For instance, ‘Hours of TV in a day’ with a mean $C_{a}$ value of 1.85 gives an $\hat{\alpha}$ of +0.92. But in many examples like ‘loads of laundry’ with mean $C_{a}$ of 2.06 gives an $\hat{\alpha}$ of +1.87. $\hat{\alpha}>1$ implies the sample not only deviates to the ideal side of average but goes beyond ideal in the same direction.
We also perform the experiment with temperature zero. While higher temperatures give greater degrees of randomness to an LLM output by picking from top-k tokens, a temperature of zero means that it will always select the most probable token. With temperature zero, the $\hat{\alpha}$ of the categories remain similar ($p$ = 0.035). For instance, for the category ‘Hours of TV in a day’, with temperature zero, we get $C_{a}=3.5$, $C_{i}=2$ and $C_{s}=2.5$. The $\hat{\alpha}$ is $+0.66$. This example is illustrated in Figure 1. The table of results for the run with temperature zero is in Appendix 1.2.
@@ 4.2 Evaluating value bias learnt in context
In this experiment, we evaluate the value bias of LLM for a distribution and value system learnt in context. In other words, the experiment shows that LLMs can give value bias zero-shot with newly learned distribution and values! For this, we introduce a category: a fictitious hobby ‘glubbing’. To ensure the token is not already associated with a distribution or value, we prompt the LLM with: “Suppose there is a hobby called glubbing. What is the average number of hours a person does glubbing in a week?”. The value returned is zero, and the grade, when queried for the associated grade, gives C+ (which is the middle grade on a scale of A+ to D-).
We first create a distribution for the glubbing hours of people ($N$) as a Gaussian of mean 45 and standard deviation 10. We repeat the experiment with a bi-modal Gaussian distribution with modes at 35 and 65 and a standard deviation of 5. The implementation and analysis of the two experiments are the same.
With each prompt, we give 100 samples samples from $N$ as the glubbing hours of people. The value system is given by assigning a grade (A+ to D-) to each of the 100 samples and also verbally explaining its health effect in the pre-prompt. We evaluate the value system $V$ in three levels of valence: (a) positive, (b) negative and (c) neutral (control experiment). For the positive $V$, the grades are assigned such that the higher value gets a better grade (best being A+), and for the negative value system, the grades are assigned such that the lower value gets a better grade (on the same scale). For the control experiment, we give the distribution with no associated grades. A sample positive prompt is given below:
@@
The ‘…’ corresponds to the rest of the values and grades (the prompt has a hundred samples and corresponding grades). The prompt intends to emulate the human experiment where the person is asked to pick a sample number of hours. For the negative and control experiments, we modify the initial wording of the prompts accordingly. The full prompt set is given in the Appendix 1.4. We run the experiment for positive, negative and control settings a hundred times each. This is the $C_{s}$ value for three different $V$.
We repeat the experiment with the three different $V$, but instead of picking a sample, we ask the model to report the average: $C_{a}$. The reported average is close to the true distribution average: mean($N$). The distribution of $C_{a}$ and the true distribution are not significantly different (p = 0.62,0.30 and 0.46 for the positive, negative and neutral valence under Mann-Whitney U-test).
Figure 3 shows the result of the hundred runs for the uni-modal and bi-modal distributions. In both cases, we can see that when the value system $V$ is positive, the mean of samples is higher than the mean of reported average. For the uni-modal true distribution, the mean $C_{s}$ for negative $V$ is 36.5 while the mean $C_{s}$ for positive $V$ is 46.7. The results show a strong value bias of the LLM in picking a sample.
We use the Mann-Whitney U-test to see the statistical significance of the distribution shift caused by value bias: the distribution of generated samples $C_{s}$ and the distribution of the average reported by LLM ($C_{a}$). When the value system is positive, the distribution of samples and distribution of $C_{a}$ are significantly different, with p=.$=.003$, under Mann-Whitney U-test. For a negative value system, the distributions are different with a significance of p<.$<.001$.
@@
Figure 3: Estimates of the average amount of glubbing (green) and mean of samples (red) for the unimodal (left) and bimodal (right) conditions from experiment evaluating value bias in context. The true average (mean of input distribution $N$) amounts of glubbing presented (dashed red lines) are also shown.
4.3 Value-bias in prototypes
We evaluate value-bias in prototypes across eight categories as listed in Table 2. The categories and exemplars are taken from  Bear & Knobe (2017b). For each category, we use six exemplars, which are short descriptions of items of that category. For instance, for the category ‘High-school teacher’, the first exemplar is as follows:

### 21 Feb 2024 v2

@@ 1 Introduction
Large language models (LLMs) are being deployed across a wide range of applications (Brown et al., 2020; Schick & Schütze, 2021) with an increasing socio-economic impact (Eloundou et al., 2023). The quality of responses generated by LLMs is an essential area of research from both security (Yao et al., 2023b) and utility perspectives (Achiam et al., 2023). LLMs have been found to be capable of doing mathematical reasoning (Imani et al., 2023), pragmatics (Lipkin et al., 2023), and creating high-level plans (Song et al., 2023) to varying degrees even without specific deliberation methods like CoT (Wei et al., 2022). Understanding this non-deliberate mechanism of response sampling is crucial to understanding its performance and biases.
Previous work evaluating the biases of LLMs has shown them to have significant gender bias (Kotek et al., 2023), social stereotypes (Kotek et al., 2023), and popularity bias (Hou et al., 2023). Studies specific to critical applications have also been conducted, like medical domain (Omiye et al., 2023), code generation (Huang et al., 2023), and biases specific to cultures (Khandelwal et al., 2023) or geographic regions (Manvi et al., 2024). These are relevant studies for real-world applications but often have context-based confounders. It is thus essential to focus on understanding the underlying foundational biases manifesting in different forms of biases.
In this work, we study the sampling of an LLM in light of value bias and show that it tends to favour high-value options across different scenarios. Since the underlying auto-regressive mechanism is not goal-driven, it is not trivial how, from a large number of possibilities of samples, the LLM’s output is based on value. This value bias is a tendency for the sample to deviate from the (statistically) highly likely value towards a notion of ‘ideal value’. This unexpected bias can manifest as different biases based on the value system and the evaluated context.
@@
We show that value bias manifests in cases where the notion of ‘ideal’ stems from an implicit value system of the LLM (learnt in pre-training) or a value system learnt in context. We motivate our experiments from the ideas presented by Phillips et al. (2019), which states that possibility sampling comes from a mixture of likelihood and value. We use psychology-inspired experiments for value-bias learnt in context (Bear et al., 2020), implicit value-bias (Bear et al., 2020), and show the manifestation of value bias in forming normality and prototypes (Bear & Knobe, 2017a). Importantly, the notion of ‘ideal’ can manifest in other applications like text summarization, reducing response quality.
Value bias can be useful in many applications, if the value system of the LLM aligns with the use case. Under the categories evaluated, our experiment shows a lack of correlation between the value system of humans and LLMs. Our result shows how LLMs can learn a value system in context with a single prompt with a statistically significant shift of the sample towards the ideal side of the average, as illustrated in Figure 1.
@@
  • We show that the LLM has an implicit value bias for concepts. We also show that value bias manifests even for new concepts learnt in context (zero shot).
  • We show the practical implications of value bias in LLMs’ judgment of concept prototypicality. We also show that the value bias in LLMs might not align with the value system of humans.
@@ 2.1 LLM Biases
Previous work on LLMs predominantly evaluates biases with respect to social concepts like gender, race, and popularity. There have also been investigations of biases in aspects like language style and lexical content (Wan et al., 2023). Gallegos et al. (2023) comprehensively surveys these works and presents a taxonomy of biases. This taxonomy aligns with how humans attribute meanings to these biases and their impact on society. The biases of LLMs have also been studied in the context of specific fields and applications like health care (Omiye et al., 2023; Zack et al., 2024; Thirunavukarasu et al., 2023; Hastings, 2024). These studies, however, do not go beyond the human taxonomy of biases to explore fundamental biases that, in turn, manifest in more generic real-world applications.
@@ 2.2 Understanding Heuristics in Response Sampling
Human reasoning has been broadly characterised as a two-step process of quick ‘System-1’ reasoning and a more deliberate ‘System-2’ reasoning (Daniel, 2017). Interestingly, LLMs have been conceptually compared to System1 thinking due to their automatic and quick nature (Yao et al., 2023a). In fact, recent work shows some overlaps in errors made by LLMs and humans in System1 reasoning tasks (Dasgupta et al., 2022).
Simon (1996) uses the notion of heuristics in the decision-making of ‘System-1’ mechanisms. They show the utility of ‘mental shortcuts’ to navigate countless possibilities of the search problem (Newell et al., 1972). Exploring the heuristics of LLMs can help understand the characteristics of information processing in them (Hazra et al., 2023; Shah et al., 2023; Suri et al., 2023). Therefore, the mechanism by which LLMs sample their response and the heuristics followed require further investigation.
System1 outputs play a considerable role in System2 processing because the latter depends on the former as a prior for decision making. For instance, in AlphaGo (Silver et al., 2016), the deliberate system, i.e., the MCTS algorithm, relies on the estimates of the neural network to cut down the search space of solutions. Similarly, the sampling of the LLM becomes critical for the deliberate problem solver to narrow down the space of possibilities. For example, in ToT (Yao et al., 2023a), when the LLM is combined with a symbolic solver, there is an inherent assumption that LLMs would give a good prior for the symbolic system to solve for it to execute accurate deliberation. In summary, even for problem solving at ‘System-2’, understanding sampling heuristics is of much significance.
@@ 2.3 Possibility Sampling in Humans
A growing body of literature addresses how humans and animals navigate the search problem of countless possibilities (Phillips et al., 2019; Phillips & Cushman, 2017; Mattar & Lengyel, 2022; Ross et al., 2023). In this line of work, human decision-making is thought of as a two-step process where model-free reinforcement learning is looked at as a cheap and efficient way to construct a consideration set of possible options and spend computational resources in a model-based framework to arrive at the final choice among the limited sampled candidates (Gershman et al., 2014). Increasing evidence shows the significance of likelihood and value in how humans sample possible solutions from the large space of possibilities (Bear et al., 2020; Phillips et al., 2019; Bear & Knobe, 2017b).
In short, what comes to the human mind or what it considers normal is sampled from a probability distribution of both value and probability (frequency) (Bear et al., 2020). This dual nature of thought is hypothesised to have come due to humans being goal-driven agents and engaging in value maximisation (Bear & Knobe, 2017b). In spatial navigation, experiments show that an optimal reinforcement learning agent which has the ability to recall past experiences, ordered on utility, optimises on two dimensions, namely, “gain” and “need”, where “gain” roughly refers to the value aspect of experience while “need” refers to the statistical occurrence of experience (Mattar & Daw, 2018).
Other efforts try to encode the possibility and value dimensions into the modalities ‘could’ and ‘should’ where ‘could’ shows how humans imagine possible alternatives, and ‘should’ shows the high-value possibilities. Poulsen & DeDeo (2023) show the influence of modal value when the LLM considers possibilities.
Figure 2: The figure shows the average,ideal, and sample values reported by LLM for four different categories. $\alpha$ shows the deviation in the direction of the ideal. $\hat{\alpha}$ is the normalized value of alpha, purple color shows the positive value bias. The sample of an LLM often drifts away from the average value towards an ideal value.
@@ 3.1 Implicit Bias
We evaluate the likelihood of samples and the notion of the ideal on different categories ($C$). The average value reported by the LLM for a category $C$ is ${C}_{a}$ and the ideal value is ${C}_{i}$. The sample generated by the LLM for the category is $C_{s}$. We use a variable $\alpha$ to quantify the degree of value bias. For a positive value bias, the $C_{s}$ deviate from the $C_{a}$ in the direction of $C_{i}$. We compute this direction as the positive direction of alpha.
@@
The sign of alpha is given by the relative positions of $C_{a}$, $C_{i}$, and $C_{s}$ such that $C_{s}$ is in the direction of $C_{i}$ as measured from $C_{a}$. Figure 2 shows a few different configurations of $C_{a}$, $C_{i}$, and $C_{s}$ and the corresponding $\alpha$.
We also compute $\alpha$ on a normalized scale such that $C_{a}$ is at the origin and $C_{i}$ is at unit distance from the origin in the positive direction. We call this $\hat{\alpha}$, and it enables comparison across categories with less dependency on the scale of values. It also allows comparison with values obtained in human experiments. Value bias is significant with a higher positive value of $\hat{\alpha}$. We compute $\hat{\alpha}$ as
@@ 3.2 Evaluating Value Bias for In-Context Value and Distributions
For a new category $C$ and the corresponding $N$ and $V$ introduced in context, we ask the LLM to report the average of the distribution $C_{a}$ and a sample from the distribution $C_{s}$. Note that the notion of ideal comes from $V$. To test the significance of value bias, we use the Mann-Whitney-U test to compare $C_{a}$ to $N$ and $C_{a}$ to $C_{s}$ given $V$.
@@ 3.3 Value Bias in Evaluating Prototypes
To evaluate the potential effect of value bias in applications, we assess their performance in evaluating prototypes. Prototypes are a means of understanding how people categorize objects and concepts. They are often considered the “best example” or an “average” representation of a category. For instance, when thinking of a category like “birds”, the prototype would be a common bird like a sparrow or a robin, which most closely matches the average characteristics of the category (Smith & Medin, 1981).
We evaluate if the judgment of the LLM on the prototype is driven solely by the statistical likelihood or biased by the value component. We provide a category $C$ and a corresponding exemplar. We ask the LLM to judge on three dimensions, namely the average ($C_{a}$), ideal ($C_{i}$), and prototypicality ($C_{p}$) of the exemplar. We test for significance with a binomial test if the sampled prototypicality $C_{p}$ is pushed away from $C_{a}$ in the direction of $C_{i}$.
@@ 4 Experiments and Results
In this section, we present the three experiments for showing value bias: (a) implicit with each concept, (b) learnt in context and (c) in prototypes of concepts. Our results show a significant value bias in the sampling of LLMs. Unless mentioned otherwise, we use GPT-4 for all our experiments. Also, all prompts were appended with a pre-prompt to get floating-point numbers as answers. The complete text used in the prompts for each experiment is given in the Appendix.
@@ 4.1 Implicit Value Bias
We use a third prompt to get a sample $C_{s}$ from each category. To get a value of $C_{s}$, we use a prompt similar to the question used in human studies (Bear et al., 2020). For instance, the prompt for the category ‘TV watching of people’ is as follows:
$P_{s}$: Enter the first number that comes to mind when answering the question. Note that there is no correct answer. Please be spontaneous in the judgment: What is the number of hours of TV for a person to watch in a day?
The value reported by the LLM is averaged over ten runs reported in Table 1. The entries where $\alpha$ is positive are shown in purple colour. These categories show a value bias in sampling.
@@
Table 1: The table shows the average, ideal and sample values for the 36 different categories in the implicit bias experiment. The categories with a positive $\alpha$ (showing a value bias) are shown in purple colour.
Of the 1200 runs, 21 experiments were rejected due to the LLM not generating scores for the prompt. 1200 runs correspond to the ten iterations of the three prompts across 40 categories: average, ideal, and sample. We hypothesize that the LLM might need a jailbreak to give output to queries like: “Ideal amount of parking tickets a person should get in a year”. We analyze the 36 categories where we get consistent results.
Results: From Table 1, we can see that 24 out of 36 categories are on the ideal side of the average when sampled ($p$ = 0.033), which is statistically significant, indicating a value bias in sampling. The sample is expected to lie in between the average and ideal due to the value bias. For instance, ‘Hours of TV in a day’ with a mean $C_{a}$ value of 1.85 gives an $\hat{\alpha}$ of +0.92. But in many examples like ‘loads of laundry’ with mean $C_{a}$ of 2.06 gives an $\hat{\alpha}$ of +1.87. $\hat{\alpha}>1$ implies the sample not only deviates to the ideal side of average but goes beyond ideal in the same direction.
We also perform the experiment with temperature zero. While higher temperatures give greater degrees of randomness to an LLM’s output by picking from top-k tokens, a temperature of zero means that it will always select the most probable token. With temperature zero, the $\hat{\alpha}$ of the categories remain similar ($p$ = 0.035). For instance, for the category ‘Hours of TV in a day’, with temperature zero, we get $C_{a}=3.5$, $C_{i}=2$, and $C_{s}=2.5$. The $\hat{\alpha}$ is $+0.66$. This example is illustrated in Figure 1. The table of results for the run with temperature zero is in Appendix 1.2.
@@ 4.2 Evaluating Value Bias Learnt In Context
In this experiment, we evaluate the value bias of an LLM for a distribution and value system learnt in context. In other words, the experiment shows that LLMs can give value bias in a zero-shot way with newly learned distribution and values. For this, we introduce a category: a fictitious hobby ‘glubbing’. To ensure the token is not already associated with a distribution or value, we prompt the LLM with: “Suppose there is a hobby called glubbing. What is the average number of hours a person does glubbing in a week?”. The value returned is zero, and the grade, when queried for the associated grade, gives C+ (which is the middle grade on a scale of A+ to D-).
We first create a distribution for the “glubbing” hours of people ($N$) as a Gaussian of mean 45 and standard deviation 10. We repeat the experiment with a bi-modal Gaussian distribution with modes at 35 and 65 and a standard deviation of 5. The implementation and analysis of the two experiments are the same.
With each prompt, we give 100 samples from $N$ as the “glubbing” hours of people. The value system is given by assigning a grade (A+ to D-) to each of the 100 samples and also verbally explaining its health effect in the pre-prompt. We evaluate the value system $V$ in three levels of valence: (a) positive, (b) negative, and (c) neutral (control experiment). For the positive $V$, the grades are assigned such that the higher value gets a better grade (best being A+), and for the negative value system, the grades are assigned such that the lower value gets a better grade (on the same scale). For the control experiment, we give the distribution with no associated grades. A sample positive prompt is given below:
@@
The ‘…’ corresponds to the rest of the values and grades (the prompt has a hundred samples and corresponding grades). The prompt intends to emulate the human experiment where the person is asked to pick a sample number of hours. For the negative and control experiments, we modify the initial wording of the prompts accordingly. The full prompt set is given in the Appendix 1.4. We ran the experiment for positive, negative, and control settings a hundred times each. This is the $C_{s}$ value for three different $V$.
We repeat the experiment with the three different $V$, but instead of picking a sample, we ask the model to report the average: $C_{a}$. The reported average is close to the true distribution average: mean($N$). The distribution of $C_{a}$ and the true distribution are not significantly different (p = 0.62, 0.30, and 0.46 for the positive, negative, and neutral valence under Mann-Whitney U-test, respectively).
Results: Figure 3 shows the result of the hundred runs for the uni-modal and bi-modal distributions. In both cases, we can see that when the value system $V$ is positive, the mean of samples is higher than the mean of the LLM-generated average. For the uni-modal true distribution, the mean $C_{s}$ for negative $V$ is 36.5 while the mean $C_{s}$ for positive $V$ is 46.7. The results show a strong value bias of the LLM in picking a sample.
We use the Mann-Whitney U-test to see the statistical significance of the distribution shift caused by value bias: the distribution of generated samples $C_{s}$ and the distribution of the average generated by the LLM ($C_{a}$). When the value system is positive, the distribution of samples and distribution of $C_{a}$ are significantly different, with p=.$=.003$, under Mann-Whitney U-test. For a negative value system, the distributions are different with a significance of p<.$<.001$.
@@
Figure 3: Estimates of the average amount of glubbing (green) and mean of samples (red) for the unimodal (left) and bimodal (right) conditions from the experiment evaluating value bias in context. The true average (mean of input distribution $N$) amounts of glubbing presented (dashed red lines) are also shown.
4.3 Value Bias in Prototypes
We evaluate value bias in prototypes across eight categories as listed in Table 2. The categories and exemplars are taken from Bear & Knobe (2017b). For each category, we use six exemplars, which are short descriptions of items of that category. For instance, for the category ‘High-school teacher’, the first exemplar is as follows: